What is the point of intersection/whether the point is a solution for:
2x + y = 3
-x + 3y = -12
Thanks. x

Answer by stanbon(75887)

2x + y = 3
-x + 3y = -12
------
Modify for elimination::
2x +y = 3
-2x + 6y = -24
------------------
Add and solve for "y":
7y = -21
y = -3
------
Solve for "x":
2x + y = 3
2x -3 = 3
2x = 6
x = 3
----
Solution: (3,-3)
==========
Cheers,
Stan H.
==================